# StoxLSTM: A Stochastic Extended Long Short-Term Memory Network for Time Series Forecasting

### v1

@@ Abstract
The Extended Long Short-Term Memory (xLSTM) network has attracted widespread research interest due to its enhanced capability to model complex temporal dependencies in diverse time series applications. Despite its success, there is still potential to further improve its representational capacity and forecasting performance, particularly on challenging real-world datasets with unknown, intricate, and hierarchical dynamics. In this work, we propose a stochastic xLSTM, termed StoxLSTM, that improves the original architecture into a state space modeling framework by incorporating stochastic latent variables within xLSTM. StoxLSTM models the latent dynamic evolution through specially designed recurrent blocks, enabling it to effectively capture the underlying temporal patterns and dependencies. Extensive experiments on publicly available benchmark datasets from multiple research communities demonstrate that StoxLSTM consistently outperforms state-of-the-art baselines with better robustness and stronger generalization ability.
@@ 1 Introduction
Time series forecasting plays a vital role in numerous real-world applications, including financial market analysis [1, 2], energy systems [3, 4], weather forecasting [5], and healthcare monitoring [6]. Traditional recurrent neural networks (RNNs) [7] and their variants [8, 9], such as Long Short-Term Memory (LSTM) networks [8], have been widely used due to their capacity to model sequential data [10, 11, 12]. However, these models often struggle to capture long-range dependencies and intricate temporal structures present in many real-world time series [13].
The Extended Long Short-Term Memory network (xLSTM) [14] has emerged as a powerful architecture that enhances the conventional LSTM by introducing exponential gating mechanisms and modified memory structures. These improvements enable xLSTM to better capture complex temporal dependencies and maintain stability over long sequences. As a result, xLSTM has demonstrated superior performance in various time series tasks compared to traditional recurrent models [15, 16, 17]. Despite these advances, xLSTM remains fundamentally deterministic, limiting its ability to explicitly represent the latent temporal dynamics and stochasticity inherent in many natural processes. This limitation restricts the model’s expressiveness when dealing with highly nonlinear and uncertain time series data.
Figure 1: The recurrent unit of StoxLSTM. Diamonds denote deterministic variables, while circles denote stochastic variables. (a) shows the stochasticized sLSTM block, while (b) shows the stochasticized mLSTM block. $z_{t}$ represents the latent state obtained by stochastic sampling through the reparameterization process.
To overcome these challenges, we propose a stochastic xLSTM termed StoxLSTM by integrating xLSTM within a specially designed stochastic state space modeling (SSM) framework [18]. Rather than treating RNN and SSM as separate components to be combined [19, 20], our approach integrates latent variable dynamics directly into the recurrent structure, effectively reinterpreting the temporal evolution as a stochastic process parameterized by enhanced xLSTM blocks, as shown in Fig. 1. This formulation allows the model to capture complex hidden temporal patterns and uncertainties that go beyond the capabilities of xLSTM networks and their deterministic variants. By integrating stochastic latent dynamics directly into the recurrent architecture, our approach enhances the representational capacity beyond conventional deterministic frameworks, allowing for more flexible and robust modeling of complex temporal behaviors.
The superiority of the proposed StoxLSTM is verified through extensive experiments across a wide range of publicly available benchmark datasets covering diverse domains such as Electricity, Weather, ETT, Solar, Traffic, PEMS, and ILI [21, 22]. The experimental results demonstrate that StoxLSTM consistently outperforms state-of-the-art baselines across multiple evaluation metrics with better robustness and stronger generalization ability. These results validate the superiority and effectiveness of our stochastic state space reformulation in advancing the modeling power of recurrent architectures for time series forecasting. The main contribution of this work can be concluded as follows:
  • To the best of our knowledge, a stochastic xLSTM, termed StoxLSTM, is proposed for the first time. The StoxLSTM improves the raw deterministic xLSTM architecture by integrating stochastic latent variables into the dynamic recurrent evolutions.
  • Efficient parameterization methods of latent temporal dynamics in StoxLSTM is designed. Through the stochasticized xLSTM blocks, the StoxLSTM is enabled to flexibly capture complex, nonlinear, and stochastic dependencies that inherently exists in many time series analysis applications.
  • Comprehensive experiments on a variety of public avaliable benchmark datasets are conducted. Experimental results verified that StoxLSTM can achieve consistent and significant performance gain over current state-of-the-art methods. The superiority and effectiveness of the stochastic state space reformulation of xLSTM are validated.
@@ 2.1 Time Series Forecasting Models
Time series forecasting aims to predict future values based on historical observations, and has seen significant advances through various modeling paradigms. Among these, Transformer-based models [23] have attracted extensive attention due to their powerful sequence modeling capabilities. Notable variants include Informer [13], Autoformer [24], Crossformer [25], and FEDformer [26], which improve efficiency and accuracy for long-term TSF. However, recent work such as LTSF-Linear [27] challenges the supremacy of Transformers by demonstrating that simple linear models can outperform them on long-term TSF. These linear models [27, 28, 29] better preserve temporal order, effectively input longer look-back lengths without overfitting, and benefit from improved embedding strategies, leading to enhanced interpretability [30]. Inspired by these advances, models like iTransformer [31] and PatchTST [32] incorporate linear transformations along the temporal dimension to enhance Transformer-based forecasting.
In parallel, recurrent architectures remain competitive. xLSTM [14] has recently gained prominence due to its ability to capture complex temporal dependencies. Variants such as xLSTMTime [17], P-sLSTM [15], and xLSTM-Mixer [16] have demonstrated superior TSF performance compared to both Transformer- and linear-based models, highlighting the effectiveness of xLSTM in modeling long-range temporal patterns.
@@ 2.2 Generative Modeling for Time Series
Generative modeling approaches have emerged as powerful tools for time series analysis, particularly in capturing uncertainty and complex temporal dynamics beyond deterministic predictions. These methods often introduce latent variables or probabilistic frameworks to model the stochastic nature of real-world time series. Classical statistical models such as ARIMA [33], VAR [34] and SSMs [18] have long been used for probabilistic forecasting by modeling trends, seasonality, and inter-variable dependencies. However, these methods often rely on strong assumptions and are limited in their ability to handle nonlinearities and high-dimensional data commonly found in real-world applications.
Recent advances in deep generative modeling have introduced more flexible frameworks that can better capture complex data distributions without restrictive parametric assumptions. Generative adversarial networks (GANs) [35, 36] and diffusion-based models [37, 38] have demonstrated strong capabilities in modeling intricate temporal patterns and generating realistic synthetic sequences. Among these, deep SSMs [19, 11, 39] stand out by explicitly incorporating latent variables to represent hidden states evolving over time, thereby effectively capturing stochastic temporal dependencies. These deep SSMs typically leverage variational inference techniques [40] to approximate intractable posterior distributions, enabling efficient learning and inference. For instance, ProTran [39] combines SSMs with Transformer architectures to model long-term dependencies in a probabilistic manner, achieving strong performance in uncertainty-aware forecasting tasks.
@@ 3.1 Modeling Time series Forecasting with a Deep SSM
Consider a multivariate time series $x_{1:L+T}^{C}$, where $C$ denotes the number of variables (dimensions), $L$ denotes the look-back length, and $T$ denotes the prediction horizon. Each dimension $x_{1:L+T}^{i}$ can be treated as an individual univariate time series.
For the univariate time series $x_{1:L+T}=\{x_{1},x_{2},\ldots,x_{L},\ldots,x_{L+T}\}$, forecasting with deep learning models involves parameterizing the predictive distribution $p_{\theta}(x_{L+1:L+T}|x_{1:L})$ with network parameters $\theta$. When integrating deep learning with SSMs, latent states $z_{1:L+T}$ are introduced to capture latent dynamics, transforming the predictive distribution into the joint distribution ${p_{\theta}}\left({{x_{1:L+T}},{z_{1:L+T}}|{x_{1:L}}}\right)$.
@@ 3.1.1 Dependency Mapping of the Designed SSM
To enable efficient computation and reduce error propagation during forecasting, the generative and inference models in StoxLSTM are designed following a stochastic SSM structure with explicit dependency assumptions, as illustrated in Fig. 2. Specifically, the latent state $z_{t}$ is modeled as a Markov process, capturing the temporal evolution of latent features—meaning $x_{t}$ does not depend on previously generated observations but only on the latent state $z_{t}$ and the known history $x_{1:L}$. Formally, the joint distribution factorizes as:
Figure 2: State space models corresponding to the generative model and the inference models in StoxLSTM.
@@
$\begin{array}[]{l}\quad{p_{\theta}}\left({{x_{1:L+T}},{z_{1:L+T}}|{x_{1:{\rm{L}}}}}\right)\\ =\prod\limits_{t=1}^{L+T}{{p_{\theta}}\left({{x_{t}}|{x_{1:t-1}},{z_{1:t}},{x_{1:L}}}\right)\cdot{p_{\theta}}\left({{z_{t}}|{x_{1:t-1}},{z_{1:t-1}},{x_{1:L}}}\right)}\\ =\prod\limits_{t=1}^{L}{{p_{\theta}}\left({{x_{t}}|{z_{t}},{x_{1:t-1}}}\right)\cdot{p_{\theta}}\left({{z_{t}}|{z_{t-1}},{x_{1:t-1}}}\right)}\cdot\prod\limits_{t=L+1}^{L+T}{{p_{\theta}}\left({{x_{t}}|{z_{t}},{x_{1:L}}}\right)\cdot{p_{\theta}}\left({{z_{t}}|{z_{t-1}},{x_{1:L}}}\right)}\end{array}$ (1)
@@
wherein $\prod\limits_{t=1}^{L}{{p_{\theta}}\left({{x_{t}}|{z_{t}},{x_{1:t-1}}}\right)\cdot{p_{\theta}}\left({{z_{t}}|{z_{t-1}},{x_{1:t-1}}}\right)}$ corresponds to the reconstruction of the known history $x_{1:L}$. While $\prod\limits_{t=L+1}^{L+T}{{p_{\theta}}\left({{x_{t}}|{z_{t}},{x_{1:L}}}\right)\cdot{p_{\theta}}\left({{z_{t}}|{z_{t-1}},{x_{1:L}}}\right)}$ models the evolution of future states and observations based on the known past $x_{1:L}$.
Correspondingly, the approximate posterior of latent states $z_{1:T}$ with network parameters $\varphi$ factorizes as:
${q_{\varphi}}\left({{z_{1:L+T}}|{x_{1:L+T}}}\right)=\prod\limits_{t=1}^{L+T}{{q_{\varphi}}\left({{z_{t}}|{z_{1:t-1}},{x_{1:L+T}}}\right)}=\prod\limits_{t=1}^{L+T}{{q_{\varphi}}\left({{z_{t}}|{z_{t-1}},{x_{1:L+T}}}\right)}$ (2)
3.2 A Stochastic Extend Long Short-term Memory Network
@@
The overall structure of StoxLSTM is illustrated in Fig. 3. StoxLSTM performs channel-independent forecasting for multivariate time series by dividing the multivariate time series into multiple individual univariate sequences. Each univariate sequence $x{(i)}\in\mathbb{R}^{1\times L},i=1,2,\ldots,C$ is processed independently by the StoxLSTM module, and the outputs are concatenated to form the multivariate prediction $\widehat{x}\in\mathbb{R}^{C\times T}$. To address the challenges of modeling long sequences faced by xLSTM [15], each individual sequence is divided into multiple consecutive patches.
Figure 3: Overall framework of StoxLSTM.

### v2

@@ Abstract
The Extended Long Short-Term Memory (xLSTM) network has demonstrated strong capability in modeling complex long-term dependencies in time series data. Despite its success, the deterministic architecture of xLSTM limits its representational capacity and forecasting performance, especially on challenging real-world time series datasets characterized by inherent uncertainty, stochasticity, and complex hierarchical latent dynamics. In this work, we propose StoxLSTM, a stochastic xLSTM within a designed state space modeling framework, which integrates latent stochastic variables directly into the recurrent units to effectively model deep latent temporal dynamics and uncertainty. The designed state space model follows an efficient non-autoregressive generative approach, achieving strong predictive performance without complex modifications to the original xLSTM architecture. Extensive experiments on publicly available benchmark datasets demonstrate that StoxLSTM consistently outperforms state-of-the-art baselines, achieving superior performance and generalization.
@@ 1 Introduction
Time series forecasting plays a crucial role across diverse domains such as finance [30, 25], energy [43, Uremović_New_2023], weather [4], and healthcare [27]. Recurrent neural networks (RNNs), particularly Long Short-Term Memory (LSTM) networks [17], have been widely used due to their capacity to model sequential data [8, 28, 6, 34]. The Extended Long Short-Term Memory network (xLSTM) [3] further enhances conventional LSTM by introducing exponential gating mechanisms and modified memory structures, enabling better modeling of complex temporal dependencies and improved stability over long sequences. Consequently, xLSTM has demonstrated superior performance across various time series forecasting tasks compared to traditional recurrent models [19, 20, 1].
Despite these advances, prior work has predominantly focused on architectural modifications at the top level of xLSTM, without fundamentally redesigning its recurrent units to better capture the intrinsic characteristics of time series data [1, 20, 19]. Real-world time series often exhibit deep latent dependencies [11, 13] as well as inherent uncertainty and stochasticity [30, 10, 7], which call for explicit modeling. However, existing xLSTM-based models lack explicit modeling of deep hierarchical latent states, which restricts their capacity to fully represent complex temporal evolution patterns. Moreover, these models are entirely discriminative and deterministic, limiting their ability to model the inherent uncertainty and latent stochastic dynamics commonly present in real-world time series.
To address these limitations, we propose StoxLSTM, a stochastic xLSTM within a state space modeling framework. First, we mathematically design an efficient, non-autoregressive state space model (SSM) [11] dependency tailored to the characteristics of the xLSTM recurrent units, which accurately model the temporal evolution of latent variables. Moreover, the SSM dependency facilitates precise and efficient forecasting without introducing excessive complexity to the original xLSTM architecture. Second, we directly integrate latent variables into the xLSTM recurrent units, specifically embedding stochastic latent variables within the sLSTM and mLSTM sub-blocks to form the two stochastic recurrent units—StosLSTM and StomLSTM, as illustrated in Fig. 1. Based on the designed SSM dependency, the generative and inference models of StoxLSTM are then constructed by leveraging these stochastic recurrent units. Through variational inference, the model effectively captures meaningful latent variables, allowing it to model deep temporal dynamics and uncertainty.
Figure 1: The recurrent unit of StoxLSTM. StoxLSTM integrates stochastic latent variables based on the original xLSTM to represent more complex hierarchical and stochastic characteristics in time series data. The xLSTM consists of two sub-blocks: sLSTM and mLSTM. (a) shows the stochasticized sLSTM block, referred to as StosLSTM, while (b) shows the stochasticized mLSTM block, called StomLSTM. In the figure, diamond shapes denote deterministic variables, such as $\bm{n}_{t}$, $\bm{c}_{t}$, and $\bm{h}_{t}$ from the original xLSTM. Circles indicate stochastic variables, including the latent variable $\bm{z}_{t}$ obtained through the reparameterization module, as well as the observed time series variable $x_{t}$.
Extensive experiments on multiple publicly available benchmark datasets, such as Electricity, Weather, ETT, Solar, Traffic, PEMS, and ILI [39, 31], demonstrate that StoxLSTM consistently outperforms state-of-the-art baselines in terms of forecasting performance and generalization. Our main contributions are summarized as follows:
  • A novel, efficient non-autoregressive SSM dependency is designed, tailored to the structural characteristics of xLSTM recurrent units, enabling effective latent variable modeling and enhanced forecasting performance without introducing excessive complexity to the original architecture.
  • To the best of our knowledge, we propose StoxLSTM, the first stochastic xLSTM explicitly modeling uncertainty and complex latent dynamics in time series. Its recurrent units integrate latent stochastic variables directly into xLSTM’s sub-blocks. Following the designed SSM dependency, the generative and inference models are formulated based on these stochastic recurrent units and trained via variational inference.
  • Extensive experiments on diverse publicly available benchmark datasets demonstrate that StoxLSTM consistently outperforms state-of-the-art methods, validating the effectiveness and superiority of the stochastic state space reformulation of xLSTM.
@@ 2.1 Time Series Forecasting Models
Time series forecasting aims to predict future values based on historical observations and has witnessed significant progress through diverse modeling paradigms. Among these, Transformer-based models [33] have attracted extensive attention due to their powerful sequence modeling capabilities. Notable variants include Informer [44], Autoformer [38], Crossformer [42], and FEDformer [45], which improve efficiency and accuracy for long-term TSF. However, recent work such as LTSF-Linear [41] challenges the supremacy of Transformers by demonstrating that simple linear models can outperform them on long-term TSF. These linear models [41, 36, 35] better preserve temporal order, effectively utilize longer look-back windows without overfitting, and benefit from enhanced embedding strategies, which also improve interpretability [22]. Inspired by these advances, models such as iTransformer [23] and PatchTST [26] integrate linear transformations along the temporal dimension to further boost Transformer-based forecasting performance.
In parallel, recurrent architectures remain competitive. xLSTM [3] has recently gained prominence due to its superior ability to capture complex temporal dependencies and causal relationships inherent in time series data. Compared to Transformer-based models, xLSTM-based models exhibit stronger capabilities in modeling temporal causality, which is critical for accurate forecasting. Moreover, relative to linear models, xLSTM-based models can effectively capture more intricate nonlinear dependencies within time series. Variants such as xLSTMTime [1], P-sLSTM [19], and xLSTM-Mixer [20] have demonstrated superior TSF performance compared to both Transformer- and linear-based models, highlighting the effectiveness of xLSTM in modeling long-range temporal patterns.
@@ 2.2 Generative Modeling for Time Series
Generative modeling has become an important approach for capturing uncertainty and complex temporal dynamics beyond deterministic predictions. These methods introduce latent variables or probabilistic frameworks to model the stochastic nature of real-world time series. Classical statistical models such as ARIMA [5], VAR [24], and SSMs [11] have long been used for probabilistic forecasting by modeling trends, seasonality, and inter-variable dependencies. However, their strong assumptions and limited capacity to handle nonlinearities and high-dimensional data restrict their applicability in complex real-world scenarios.
Recent advances in deep generative modeling have introduced more flexible frameworks that can better capture complex data distributions without restrictive parametric assumptions. Generative adversarial networks (GANs) [15, 40] and diffusion-based models [16, 29] have demonstrated strong capabilities in modeling intricate temporal patterns and generating realistic synthetic sequences. Among these, deep state space models (SSMs) [13, 28, 32] stand out by explicitly incorporating latent variables to represent hidden states that evolve over time. This latent variable framework enables deep SSMs to effectively capture complex temporal dependencies that are often nonlinear and stochastic in nature. By modeling the underlying hidden dynamics, deep SSMs can naturally account for the inherent randomness and uncertainty present in real-world time series data, which traditional deterministic models may fail to represent adequately. These deep SSMs typically leverage variational inference techniques [18] to approximate intractable posterior distributions, enabling efficient learning and inference. For instance, ProTran [32] combines SSMs with Transformer architectures to model long-term dependencies in a probabilistic manner, achieving strong performance in uncertainty-aware forecasting tasks.
@@ 3.1 Modeling Time series Forecasting with a Deep SSM
Consider a multivariate time series $\bm{x}_{1:L+T}^{C}$, where $C$ denotes the number of variables (dimensions), $L$ denotes the look-back length, and $T$ denotes the prediction horizon. Each dimension $x_{1:L+T}^{i}$ can be treated as an individual univariate time series.
For the univariate time series $x_{1:L+T}=\{x_{1},x_{2},\ldots,x_{L},\ldots,x_{L+T}\}$, forecasting with deep learning models involves parameterizing the predictive distribution $p_{\bm{\theta}}(x_{L+1:L+T}|x_{1:L})$ with network parameters $\bm{\theta}$. When integrating deep learning with SSMs, latent states $\bm{z}_{1:L+T}$ are introduced to capture latent dynamics, transforming the predictive distribution into the joint distribution ${p_{\bm{\theta}}}\left({{x_{1:L+T}},{\bm{z}_{1:L+T}}|{x_{1:L}}}\right)$.
@@ 3.1.1 Dependency Mapping of the Designed SSM
Considering that the hidden state of xLSTM at each time step already encodes comprehensive information from all past observations [3], and to avoid introducing excessive complexity into the StoxLSTM recurrent units, we design the generative and inference models in StoxLSTM based on a stochastic SSM framework with explicit dependency assumptions. This design also effectively addresses the error accumulation commonly observed in multi-step autoregressive forecasting [9, 41]. The overall architecture is illustrated in Fig. 2.
Figure 2: State space models corresponding to the generative model and the inference models in StoxLSTM. (a) and (b) illustrate the reconstruction and forecasting phase of the generative model, respectively. In (a), the state space model depicts the observation $x_{t}$ following the conditional distribution $p_{\bm{\theta}}(x_{t}\mid\bm{z}_{t},x_{1:t-1})$, while the latent state $\bm{z}_{t}$ evolves according to $p_{\bm{\theta}}(\bm{z}_{t}\mid\bm{z}_{t-1},x_{1:t-1})$. (b) represents the forecasting phase, where $x_{t}$ is generated from $p_{\bm{\theta}}(x_{t}\mid\bm{z}_{t},x_{1:L})$, and the latent state $\bm{z}_{t}$ transitions according to $p_{\bm{\theta}}(\bm{z}_{t}\mid\bm{z}_{t-1},x_{1:L})$. (c) illustrates the inference model, in which the latent state $\bm{z}_{t}$ is inferred from the approximate posterior distribution $q_{\phi}(\bm{z}_{t}\mid\bm{z}_{t-1},x_{1:L+T})$.
Specifically, as shown in Fig. 2(a), the current latent variable $\bm{z_{t}}$ is designed to capture all historical observations $x_{1:t-1}$, while the evolution of latent variables follows the Markov assumption. The current observation $x_{t}$ depends exclusively on the current latent state $\bm{z_{t}}$ and the entire history of observations $x_{1:t-1}$. Notably, the SSM dependency structure differs between the reconstruction and forecasting phases: during forecasting, as shown in Fig. 2(b), the model conditions solely on the known observations $x_{1:L}$, independent of previously predicted outputs. This non-autoregressive formulation effectively mitigates error propagation in multi-step forecasting. Formally, the joint distribution factorizes as:
@@
$\begin{array}[]{l}\quad{p_{\bm{\theta}}}\left({{x_{1:L+T}},{\bm{z}_{1:L+T}}|{x_{1:{\rm{L}}}}}\right)\\ =\prod\limits_{t=1}^{L+T}{{p_{\bm{\theta}}}\left({{x_{t}}|{x_{1:t-1}},{\bm{z}_{1:t}},{x_{1:L}}}\right)\cdot{p_{\bm{\theta}}}\left({{\bm{z}_{t}}|{x_{1:t-1}},{\bm{z}_{1:t-1}},{x_{1:L}}}\right)}\\ =\prod\limits_{t=1}^{L}{{p_{\bm{\theta}}}\left({{x_{t}}|{\bm{z}_{t}},{x_{1:t-1}}}\right)\cdot{p_{\bm{\theta}}}\left({{\bm{z}_{t}}|{\bm{z}_{t-1}},{x_{1:t-1}}}\right)}\cdot\prod\limits_{t=L+1}^{L+T}{{p_{\bm{\theta}}}\left({{x_{t}}|{\bm{z}_{t}},{x_{1:L}}}\right)\cdot{p_{\bm{\theta}}}\left({{\bm{z}_{t}}|{\bm{z}_{t-1}},{x_{1:L}}}\right)}\end{array}$ (1)
@@
wherein $\prod\limits_{t=1}^{L}{{p_{\bm{\theta}}}\left({{x_{t}}|{\bm{z}_{t}},{x_{1:t-1}}}\right)\cdot{p_{\bm{\theta}}}\left({{\bm{z}_{t}}|{\bm{z}_{t-1}},{x_{1:t-1}}}\right)}$ corresponds to the reconstruction of the historical observations $x_{1:t-1}$. While $\prod\limits_{t=L+1}^{L+T}{{p_{\bm{\theta}}}\left({{x_{t}}|{\bm{z}_{t}},{x_{1:L}}}\right)\cdot{p_{\bm{\theta}}}\left({{\bm{z}_{t}}|{\bm{z}_{t-1}},{x_{1:L}}}\right)}$ models the evolution of future latent states and observations conditioned on the known observations $x_{1:L}$.
The inference model estimates the approximate posterior distribution of the latent variables. As illustrated in Fig. 2(c), the current latent variable $\bm{z_{t}}$ in the inference model has access to the entire observation $x_{1:L+T}$. The evolution of latent variables similarly follows the Markov assumption. Accordingly, the approximate posterior over the latent states $\bm{z}_{1:L+T}$, parameterized by the inference model parameters $\bm{\varphi}$, factorizes as:
${q_{\bm{\varphi}}}\left({{\bm{z}_{1:L+T}}|{x_{1:L+T}}}\right)=\prod\limits_{t=1}^{L+T}{{q_{\bm{\varphi}}}\left({{\bm{z}_{t}}|{\bm{z}_{1:t-1}},{x_{1:L+T}}}\right)}=\prod\limits_{t=1}^{L+T}{{q_{\bm{\varphi}}}\left({{\bm{z}_{t}}|{\bm{z}_{t-1}},{x_{1:L+T}}}\right)}$ (2)
3.2 A Stochastic Extended Long Short-term Memory Network
@@
The overall structure of StoxLSTM is illustrated in Fig. 3. StoxLSTM performs channel-independent forecasting for multivariate time series by dividing the multivariate time series into multiple individual univariate sequences. Each univariate sequence $\bm{x_{i}}\in\mathbb{R}^{1\times L},i=1,2,\ldots,C$ is processed independently by the StoxLSTM module, and the outputs are concatenated to form the multivariate prediction $\bm{\widehat{x}}\in\mathbb{R}^{C\times T}$. To address the challenges of modeling long sequences encountered by xLSTM [19], each univariate sequence is further segmented into consecutive patches. Section 3.2.1 introduces the two recurrent units of StoxLSTM, namely StosLSTM and StomLSTM, while Sections 3.2.2 and 3.2.3 describe how the generative and inference models of StoxLSTM are constructed based on these recurrent units.
Figure 3: Overall framework of StoxLSTM. (a) depicts the generative model, while (b) illustrates the inference model. In both (a) and (b), the dashed boxes represent the stacked recurrent units of StoxLSTM shown in (c). Each stacked unit in (c) corresponds to either a StomLSTM or a StosLSTM block.